# A Gauss–Bonnet formula for the renormalized area of minimal submanifolds of Poincaré–Einstein manifolds

### v1

@@ 1. Introduction
Poincaré–Einstein manifolds are a generalization of hyperbolic space which have been intensely studied in both geometry and physics. These are asymptotically hyperbolic manifolds $(X^{n},g_{+})$ with $\operatorname{Ric}(g_{+})=-(n-1)g_{+}$. Although their volume is infinite, they have a renormalized volume $\mathcal{V}$, which is an invariant when $n$ is even. Their minimal submanifolds $Y^{k}\subset X^{n}$ with prescribed boundary at infinity are also fundamental objects. Their area is infinite, but they have a renormalized area $\mathcal{A}$, which is an invariant when $k$ is even.
@@
A main ingredient in Chang, Qing, and Yang’s derivation of Equation (1.1) is a result of Alexakis [Alexakis2012], motivated by a conjecture of Deser and Schwimmer [DeserSchwimmer1993], that establishes in even dimension $n$ a decomposition of any natural scalar $I$ whose integral over compact manifolds is conformally invariant. Namely, any such scalar can be written
@@
where $c\in\mathbb{R}$, $\mathrm{Pf}$ denotes the Pfaffian of the curvature tensor, $\mathcal{Z}_{I}$ is a natural scalar that is pointwise conformally invariant of weight $-n$, and $V$ is a natural vector field. In their proof of Equation (1.1), Chang, Qing, and Yang applied Alexakis’ result with $I$ equal to Branson’s [Branson1995] critical $Q$-curvature. The proof shows that the conformal invariant in (1.1) is given by $\mathcal{Z}=\frac{(-1)^{n/2}}{(n-1)!}\mathcal{Z}_{Q}$.
Our main result is an analogue of Equation (1.1) for even-dimensional minimal submanifolds of Poincaré–Einstein manifolds, assuming a special case of a submanifold version of Alexakis’ result. We first formulate this submanifold version as a conjecture.
@@
By convention, manifolds and submanifolds are without boundary unless otherwise specified. See Section 2 for a discussion of natural submanifold tensors.
An extrinsic submanifold version of Branson’s critical $Q$-curvature was defined by Case, Graham, and Kuo [CaseGrahamKuo2023]. For $k,n\in\mathbb{N}$ with $k$ even and $n>k$, this $Q$-curvature is a natural submanifold scalar whose integral over compact $Y$ is conformally invariant. It will be reviewed in Section 2. Henceforth in this paper, by $Q$ we will always mean this critical extrinsic submanifold $Q$-curvature.
@@
Importantly, the conformal invariance of $\mathcal{W}_{Q}$ implies that $\int_{Y}\mathcal{W}_{Q}\operatorname{dA}$ converges.
@@
The main reason that it is important to have a conformally invariant integrand in Equations (1.1) and (1.4) is to guarantee convergence of the integral. The formulas then give a concrete expression for the outcome of the renormalization process applied to volume or area. There are other formulas of Gauss–Bonnet type in this setting, for instance in [Albin2009] for Poincaré–Einstein manifolds and in [TaylorToo2020] for minimal submanifolds. But these formulas involve renormalized quantities other than just the volume or area. A drawback of Equations (1.1) and (1.4) is that they require identifying a conformal invariant in the decomposition of $Q$-curvature to become explicit.
@@ 2.1. Natural submanifold tensors and extrinsic $Q$-curvature
Here $M_{j}$, $N_{j}$, and $P$ denote powers, and $\pi$ and $\pi_{j}$ denote restriction to $Y$ followed by projection to either $TY$ or $NY$ in each index. The tensor displayed at (2.1) is viewed as covariant in all indices (that is, with all indices lowered), and the contractions are taken with respect to the metrics induced by $g$ on $TY$ and $NY$ for partial pairings of tangential and normal indices. A natural submanifold tensor $T$ has weight $w\in\mathbb{R}$ if $T^{c^{2}g}=c^{w}T^{g}$ for all $c>0$. A natural submanifold tensor $T$ is conformally invariant of weight $w$, or a conformal submanifold invariant, if $T^{e^{2\Upsilon}g}=e^{wi^{\ast}\Upsilon}T^{g}$ for all immersions $i\colon Y^{k}\to(X^{n},g)$ and all $\Upsilon\in C^{\infty}(X)$.
@@
Note that $P_{k}$ and $Q$ depend on the conformal class $[g]$ on $X$ and the choice of $h\in i^{\ast}[g]$. We can also view $P_{k}$ and $Q$ as determined simply by a choice of metric $g$ on $X$, since $g$ determines $[g]$ and the induced metric $h=i^{*}g$. When viewed this way, $P_{k}$ is a natural submanifold differential operator and $Q$ is a natural submanifold scalar. (A different definition of natural submanifold differential operators and scalars is used in [CaseGrahamKuo2023]. The definitions will be shown to be equivalent in [GrahamKuo2023].)
@@
An essential property of the operators and $Q$-curvatures constructed in [CaseGrahamKuo2023] is their factorization for minimal submanifolds of Einstein manifolds. We rely on the specialization of this fact in the critical-order case:
@@ 2.2. Asymptotics and scattering compactification
Let $(X^{n},g_{+})$, $n\geq 3$, be a Poincaré–Einstein manifold with conformal infinity $(\partial X,\mathfrak{c})$. That is, $X$ is a compact connected manifold with nonempty boundary, $g_{+}$ is an asymptotically hyperbolic metric satisfying $\operatorname{Ric}(g_{+})=-(n-1)g_{+}$, and $\mathfrak{c}=\bigl{[}r^{2}g_{+}\rvert_{T\partial X}\bigr{]}$, where $r$ is a defining function for $\partial X$. A representative $g_{(0)}\in\mathfrak{c}$ uniquely determines [GrahamLee1991]*Lemma 5.2 a defining function $r$ near $\partial X$ such that $r^{2}g_{+}|_{T\partial X}=g_{(0)}$ and $\lvert dr\rvert^{2}_{r^{2}g_{+}}=1$. We call $r$ the geodesic defining function and $r^{2}g_{+}$ the geodesic compactification determined by $g_{(0)}$. The metric $g_{(0)}$ also determines, for some $\epsilon>0$, a boundary identification with respect to which
@@
We have used a ‘0’ index for the $r$-direction and commas denote partial derivatives with respect to the coordinates $(x^{\alpha},r)$ on $Y$. The components of $\overline{h}$ and the derivatives of $u$ are evaluated at $(x,r)$. The above formulas for the components of $\overline{h}$ were obtained from the pullback of $r^{2}g_{+}$ upon writing
@@
Set $h_{(0)}:=i^{*}g_{(0)}$. Let $\rho$ be the geodesic defining function for $h_{+}$ determined by $h_{(0)}$ and let $(y,\rho)$, $y\in\partial Y$, be the associated boundary identification for $Y$. For this discussion we identify $Y$ with $i(Y)$ near $\partial Y$. The map $(x,r)\to(y,\rho)$ relating the the two boundary identifications is uniquely determined by the requirement that
@@
Consider the expansion at $r=0$ of $g=(\rho/r)^{2}(dr^{2}+g_{r})$ in the $(x,u,r)$ coordinates. Using the expansion (2.4) for $g_{r}$, the expansion (2.11) for $\rho/r$, and the fact that $\rho$ was extended to be independent of $u$, it follows that the expansion of $g$ is even through order $k-2$ and has no $r^{k-1}$ term. An evenness argument implies $\langle\mathsf{n},V^{g}\rangle=0$: Define $g^{0}$ by truncating the expansion of $g$ at order $k-1$. Likewise, define $U^{0}_{r}$ by truncating the expansion (2.6) of $U_{r}$ at order $k$ (keeping the term of order $k$). Define $Y^{0}$ by (2.5) with $U_{r}$ replaced by $U^{0}_{r}$ and define $V^{0}$ to be the natural vector field determined by $g^{0}$ and $Y^{0}$. Then $V=V^{0}$ at $r=0$. Since they are polynomial and even in $r$, both $g^{0}$ and $U^{0}$ extend to $r\in(-\epsilon,\epsilon)$ and are invariant under the reflection $\mathcal{R}(p,r):=(p,-r)$ on $\partial X\times(-\epsilon,\epsilon)_{r}$. Extend $Y^{0}$ to $r<0$ by (2.5) with $Y$ replaced by $Y^{0}$ and $U_{r}$ replaced by $U_{r}^{0}$. Naturality implies that *$\mathcal{R}^{*}V^{0}=V^{0}$. Hence *$\mathcal{R}^{*}V=V$ at $r=0$. But *$\mathcal{R}^{*}\mathsf{n}=-\mathsf{n}$, and hence *$\mathcal{R}^{*}\langle\mathsf{n},V\rangle=-\langle\mathsf{n},V\rangle$ at $r=0$. We conclude that $\langle\mathsf{n},V\rangle=0$.
@@ 4.1. Conventions from Riemannian geometry
respectively. Note that our convention for the Cotton tensor differs from that used in [CaseGrahamKuo2023]. Clearly $C_{abc}=-C_{bac}$. We use square brackets to denote skew symmetrization of the enclosed indices; e.g.
$C_{[ab]c}:=\frac{1}{2}(C_{abc}-C_{bac})=C_{abc}.$
@@ 4.3. Conformal submanifolds
for all Riemannian metrics $g$ on $X$ and all $\Upsilon\in C^{\infty}(X)$. It is straightforward to check that ${}^{\bullet}$ satisfies a Leibniz rule: Suppose that $L_{1}\colon C^{\infty}(Y;T_{1})\to C^{\infty}(Y;T_{2})$ and $L_{2}\colon C^{\infty}(Y;T_{2})\to C^{\infty}(Y;T_{3})$ have homogeneities $h_{1}$ and $h_{2}$, respectively. Fix $w\in\mathbb{R}$. Then
@@ 6.2. The four-dimensional case
$I\equiv-a_{1}H^{\alpha^{\prime}}\overline{\Delta}H_{\alpha^{\prime}}-3a_{4}\lvert H\rvert^{4}-a_{5}\lvert H\rvert^{2}\lvert\mathring{L}\rvert^{2}-a_{6}H^{\alpha^{\prime}}\mathring{L}_{\alpha\beta\alpha^{\prime}}\mathring{L}^{\alpha\beta\beta^{\prime}}H_{\beta^{\prime}}\\ -a_{7}\lvert H\rvert^{2}\overline{\mathsf{J}}-a_{8}\mathsf{G}\lvert H\rvert^{2}-a_{13}H^{\alpha^{\prime}}H^{\beta^{\prime}}W_{\alpha^{\prime}\alpha\beta^{\prime}}{}^{\beta}\mod\operatorname{span}\mathscr{I}_{4}^{0}.$
Repeating this argument twice more yields $I\in\operatorname{span}\mathscr{I}_{4}^{0}$.
@@
[…]
@@
[…]

### v2

@@ 1. Introduction
Poincaré–Einstein manifolds are asymptotically hyperbolic $n$-manifolds with constant Ricci curvature $\operatorname{Ric}=-(n-1)g$; see Section 2.2 for a precise definition. The motivating example is the Poincaré ball model of hyperbolic space. They have a boundary at infinity which inherits a conformal class of metrics. Poincaré–Einstein manifolds were introduced by Fefferman and Graham [FeffermanGraham1985] as a means of studying conformal geometry by associating a single Riemannian metric to a conformal class, and have been intensely studied from that point of view. Later it was realized that their geometry is precisely that underlying the AdS/CFT correspondence in physics [Maldacena1998, GubserKlebanovPolyakov1998, Witten1998, HenningsonSkenderis1998].
The boundary problem of finding a minimal submanifold of hyperbolic space with prescribed boundary at infinity has also been much studied, beginning with work of Anderson [Anderson1982, Anderson1983]. In the AdS/CFT correspondence, such minimal submanifolds $Y^{k}$ of Poincaré–Einstein manifolds arise as dual objects to submanifold observables in the boundary theory [GrahamWitten1999].
Poincaré–Einstein $n$-manifolds have infinite volume. But they have a renormalized volume $\mathcal{V}$, which is an invariant when $n$ is even. Similarly, a minimal submanifold $Y^{k}$ with prescribed boundary at infinity has infinite area, but there is a renormalized area $\mathcal{A}$ which is an invariant when $k$ is even. The renormalized volume and renormalized area both arose first in the physics literature and they are now regarded as fundamental geometric invariants; see [Graham2000] for further references and discussion. One would like to determine the relationship between these and other geometric invariants of Poincaré–Einstein manifolds and their minimal submanifolds.
@@
A main ingredient in Chang, Qing, and Yang’s derivation of Equation (1.1) is a result of Alexakis [Alexakis2012], motivated by a conjecture of Deser and Schwimmer [DeserSchwimmer1993], that establishes a decomposition of any natural scalar whose integral over compact manifolds is conformally invariant. Namely, any such scalar $I$ on manifolds of even dimension $n$ can be written
@@
where $c\in\mathbb{R}$, $\mathrm{Pf}$ denotes the Pfaffian of the curvature tensor, $\mathcal{Z}_{I}$ is a natural scalar that is pointwise conformally invariant of weight $-n$, and $V$ is a natural vector field. Branson’s [Branson1995, ChangEastwoodOrstedYang2008] critical $Q$-curvature is a natural scalar that plays a fundamental role in conformal geometry. It is not conformally invariant, but it transforms by addition of a divergence under conformal change. Hence its integral over compact manifolds is conformally invariant. Chang, Qing, and Yang applied Alexakis’ result with $I=Q$ in their proof of Equation (1.1). The proof shows that the conformal invariant in (1.1) is given by $\mathcal{Z}=\frac{(-1)^{n/2}}{(n-1)!}\mathcal{Z}_{Q}$.
Our main result is an analogue of Equation (1.1) for even-dimensional minimal submanifolds of Poincaré–Einstein manifolds, assuming a special case of a submanifold version of Alexakis’ result. We first formulate this submanifold version as a conjecture. Our convention is that manifolds and submanifolds are without boundary unless otherwise specified. Natural submanifold scalars and tensors are defined in Section 2.1.
@@
An extrinsic submanifold version of Branson’s critical $Q$-curvature was defined by Case, Graham, and Kuo [CaseGrahamKuo2023]. For $k,n\in\mathbb{N}$ with $k$ even and $n>k$, this $Q$-curvature is a natural submanifold scalar whose integral over compact $Y$ is conformally invariant and which has a nice factorization for minimal submanifolds of Einstein manifolds. It will be reviewed in Section 2. Henceforth in this paper, by $Q$ we will always mean this critical extrinsic submanifold $Q$-curvature.
@@
Importantly, the conformal invariance of $\mathcal{W}_{Q}$ implies that $\int_{Y}\mathcal{W}_{Q}\operatorname{dA}$ converges. See Section 2.2 for the definition of polyhomogeneous minimal immersions.
@@
The main reason that it is important to have a conformally invariant integrand in Equations (1.1) and (1.4) is to ensure convergence of the integral. The formulas then give a concrete expression for the outcome of the renormalization process applied to volume or area. There are other formulas of Gauss–Bonnet type in this setting, for instance in [Albin2009] for Poincaré–Einstein manifolds and in [TaylorToo2020] for minimal submanifolds. But these formulas involve renormalized quantities other than just the volume or area. A drawback of Equations (1.1) and (1.4) is that they require identifying a conformal invariant in the decomposition of $Q$-curvature to become explicit.
@@ 2.1. Natural submanifold tensors and extrinsic $Q$-curvature
Here $M_{j}$, $N_{j}$, and $P$ denote powers, and $\pi$ and $\pi_{j}$ denote restriction to $Y$ followed by projection to either $TY$ or $NY$ in each index. The tensor displayed at (2.1) is viewed as covariant in all indices (that is, with all indices lowered), and the contractions are taken with respect to the metrics induced by $g$ on $TY$ and $NY$ for partial pairings of tangential and normal indices. The special case $r,s=0$ defines natural submanifold scalars. A natural submanifold tensor $T$ has weight $w\in\mathbb{R}$ if $T^{c^{2}g}=c^{w}T^{g}$ for all $c>0$. A natural submanifold tensor $T$ is conformally invariant of weight $w$, or a conformal submanifold invariant, if $T^{e^{2\Upsilon}g}=e^{wi^{\ast}\Upsilon}T^{g}$ for all immersions $i\colon Y^{k}\to(X^{n},g)$ and all $\Upsilon\in C^{\infty}(X)$.
@@
The paper [CaseGrahamKuo2023] uses a different definition of natural submanifold scalar differential operators, and hence of natural submanifold scalars, that is formulated in terms of isometry invariant assignments depending polynomially on the metric. Graham and Kuo extended [GrahamKuo2023] that definition to tensors and showed that it is equivalent to the definition in this paper.
@@
Note that $P_{k}$ and $Q$ depend on the conformal class $[g]$ on $X$ and the choice of $h\in i^{\ast}[g]$. We can also view $P_{k}$ and $Q$ as determined simply by a choice of metric $g$ on $X$, since $g$ determines $[g]$ and the induced metric $h=i^{*}g$. When viewed this way, $P_{k}$ is a natural submanifold differential operator and $Q$ is a natural submanifold scalar.
@@
There are many operators and related curvatures which satisfy the conclusions of Proposition 2.1; see [GoverWaldron2017, GoverWaldron2015] for a different family and the introduction of [CaseGrahamKuo2023] for a detailed comparison of these two families. An essential property of the operators and $Q$-curvatures constructed in [CaseGrahamKuo2023] is their factorization for minimal submanifolds of Einstein manifolds. We rely on the specialization of this fact in the critical-order case:
@@ 2.2. Asymptotics and scattering compactification
Let $(X^{n},g_{+})$, $n\geq 3$, be a Poincaré–Einstein manifold with conformal infinity $(\partial X,\mathfrak{c})$. That is, $X$ is a compact connected manifold with nonempty boundary, $g_{+}$ is a complete metric in the interior of $X$ for which $\operatorname{Ric}_{g_{+}}=-(n-1)g_{+}$, and $r^{2}g_{+}$ extends to a $C^{2,\alpha}$-metric on $X$ such that $\mathfrak{c}:=\bigl{[}r^{2}g_{+}\rvert_{T\partial X}\bigr{]}$ is a smooth conformal class whenever $r$ is a defining function for $\partial X$. Poincaré–Einstein manifolds are necessarily asymptotically hyperbolic; i.e. $\lvert dr\rvert_{r^{2}g_{+}}=1$ along $\partial X$ for any defining function $r$ for $\partial X$. A representative $g_{(0)}\in\mathfrak{c}$ uniquely determines [GrahamLee1991]*Lemma 5.2 a defining function $r$ near $\partial X$ such that $r^{2}g_{+}|_{T\partial X}=g_{(0)}$ and $\lvert dr\rvert^{2}_{r^{2}g_{+}}=1$. We call $r$ the geodesic defining function and $r^{2}g_{+}$ the geodesic compactification determined by $g_{(0)}$. The metric $g_{(0)}$ also determines, for some $\epsilon>0$, a boundary identification with respect to which
@@
We have used a ‘0’ index for the $r$-direction, commas denote partial derivatives with respect to the coordinates $(x^{\alpha},r)$ on $Y$, and round parentheses denote symmetrization as in (4.3). The components of $\overline{h}$ and the derivatives of $u$ are evaluated at $(x,r)$. The above formulas for the components of $\overline{h}$ were obtained from the pullback of $r^{2}g_{+}$ upon writing
@@
Set $h_{(0)}:=i^{*}g_{(0)}$. Let $\rho$ be the geodesic defining function for $h_{+}$ determined by $h_{(0)}$ and let $(y,\rho)$, $y\in\partial Y$, be the associated boundary identification for $Y$. For this discussion we identify $Y$ with $i(Y)$ near $\partial Y$. The map $(x,r)\to(y,\rho)$ relating the two boundary identifications is uniquely determined by the requirement that
@@
Consider the expansion at $r=0$ of $g=(\rho/r)^{2}(dr^{2}+g_{r})$ in the $(x,u,r)$ coordinates. Using the expansion (2.4) for $g_{r}$, the expansion (2.11) for $\rho/r$, and the fact that $\rho$ was extended to be independent of $u$, it follows that the expansion of $g$ is even through order $k-2$ and has no $r^{k-1}$ term. An evenness argument implies $\langle\mathsf{n},V^{g}\rangle=0$: Define $g^{0}$ by truncating the expansion of $g$ at order $k-1$. Likewise, define $U^{0}_{r}$ by truncating the expansion (2.6) of $U_{r}$ at order $k$ (keeping the term of order $k$). Define $Y^{0}$ by (2.5) with $U_{r}$ replaced by $U^{0}_{r}$ and define $V^{0}$ to be the natural vector field determined by $g^{0}$ and $Y^{0}$. Then $V=V^{0}$ at $r=0$. Since they are polynomial and even in $r$, both $g^{0}$ and $U^{0}$ extend to $r\in(-\epsilon,\epsilon)$ and are invariant under the reflection $\mathcal{R}(p,r):=(p,-r)$ on $\partial X\times(-\epsilon,\epsilon)_{r}$. Extend $Y^{0}$ to $r<0$ by (2.5) with $Y$ replaced by $Y^{0}$ and $U_{r}$ replaced by $U_{r}^{0}$. Naturality implies that $\mathcal{R}^{*}V^{0}=V^{0}$. Hence $\mathcal{R}^{*}V=V$ at $r=0$. But $\mathcal{R}^{*}\mathsf{n}=-\mathsf{n}$, and hence $\mathcal{R}^{*}\langle\mathsf{n},V\rangle=-\langle\mathsf{n},V\rangle$ at $r=0$. We conclude that $\langle\mathsf{n},V\rangle=0$.
@@ 4.1. Conventions from Riemannian geometry
respectively. Note that our convention for the Cotton tensor differs from that used in [CaseGrahamKuo2023]. Clearly $C_{abc}=-C_{bac}$. We use square brackets (resp. round parentheses) to denote skew symmetrization (resp. symmetrization) of the enclosed indices; e.g.
(4.3) $\begin{split}T_{[abc]}&:=\frac{1}{3!}\left(T_{abc}-T_{bac}-T_{cba}-T_{acb}+T_{bca}+T_{cab}\right),\\ T_{(abc)}&:=\frac{1}{3!}\left(T_{abc}+T_{bac}+T_{cba}+T_{acb}+T_{bca}+T_{cab}\right).\end{split}$
@@ 4.3. Conformal submanifolds
for all Riemannian metrics $g$ on $X$ and all $\Upsilon\in C^{\infty}(X)$. It is straightforward to check that satisfies a Leibniz rule: Suppose that $L_{1}\colon C^{\infty}(Y;T_{1})\to C^{\infty}(Y;T_{2})$ and $L_{2}\colon C^{\infty}(Y;T_{2})\to C^{\infty}(Y;T_{3})$ have homogeneities $h_{1}$ and $h_{2}$, respectively. Fix $w\in\mathbb{R}$. Then
@@ 6.2. The four-dimensional case
$I\equiv-a_{1}H^{\alpha^{\prime}}\overline{\Delta}H_{\alpha^{\prime}}-3a_{4}\lvert H\rvert^{4}-a_{5}\lvert H\rvert^{2}\lvert\mathring{L}\rvert^{2}-a_{6}H^{\alpha^{\prime}}\mathring{L}_{\alpha\beta\alpha^{\prime}}\mathring{L}^{\alpha\beta\beta^{\prime}}H_{\beta^{\prime}}\\ -a_{7}\lvert H\rvert^{2}\overline{\mathsf{J}}-a_{8}\mathsf{G}\lvert H\rvert^{2}-a_{13}H^{\alpha^{\prime}}H^{\beta^{\prime}}W_{\alpha^{\prime}\alpha\beta^{\prime}}{}^{\alpha}\mod\operatorname{span}\mathscr{I}_{4}^{0}.$
Again let $\Upsilon\in C^{\infty}(X)$ be such that $\Upsilon\rvert_{Y}=0$. Then
00 0 $=\int_{Y}\left(I\operatorname{dA}\right)^{\bullet}$
$=\int_{Y}\Bigl{(}2a_{1}\overline{\Delta}H_{\alpha^{\prime}}+12a_{4}\lvert H\rvert^{2}H_{\alpha^{\prime}}+2a_{5}\lvert\mathring{L}\rvert^{2}H_{\alpha^{\prime}}+2a_{6}H^{\beta^{\prime}}\mathring{L}_{\alpha\beta\beta^{\prime}}\mathring{L}^{\alpha\beta}{}_{\alpha^{\prime}}$
$\qquad+2a_{7}\overline{\mathsf{J}}H_{\alpha^{\prime}}+2a_{8}\mathsf{G}H_{\alpha^{\prime}}+2a_{13}H^{\beta^{\prime}}W_{\alpha^{\prime}\alpha\beta^{\prime}}{}^{\alpha}\Bigr{)}\Upsilon^{\alpha^{\prime}}\operatorname{dA}.$
It follows that
$I\equiv 3a_{4}\lvert H\rvert^{4}\mod\operatorname{span}\mathscr{I}_{4}^{0}.$
Repeating this variational argument one more time yields $I\in\operatorname{span}\mathscr{I}_{4}^{0}$.
@@
$\left\{\mathsf{F}_{\alpha\beta},\mathring{L}_{\alpha\beta}^{2},\lvert\mathring{L}\rvert^{2}g_{\alpha\beta}\right\}\quad\text{and}\quad\left\{\overline{\nabla}^{\alpha}(\mathring{L}^{\beta\gamma\alpha^{\prime}}W_{\alpha\beta\gamma\alpha^{\prime}})\right\}$
@@
$\left\{\mathsf{F}_{\alpha\beta},\mathring{L}_{\alpha\beta}^{2},\lvert\mathring{L}\rvert^{2}g_{\alpha\beta},\mathsf{G}g_{\alpha\beta}\right\}\quad\text{and}\quad\left\{\overline{\nabla}^{\alpha}(\mathring{L}^{\beta\gamma\alpha^{\prime}}W_{\alpha\beta\gamma\alpha^{\prime}}),\overline{\nabla}^{\alpha}(\mathring{L}_{\alpha}{}^{\beta\alpha^{\prime}}W_{\beta\gamma\alpha^{\prime}}{}^{\gamma})\right\}$